# Bankrupting DoS Attackers

### v4

@@ Abstract
Keywords:
Denial-of-service attacks Resource burning Distributed algorithms.
@@ 1 Introduction
It has long been known that fees can be used to curb DoS attacks [44, 63, 7, 117, 88, 121, 99]. This fee may be in the form of computation: a client solves a computational puzzle or Captcha per request [65, 43, 83, 62]; or bandwidth: a client sends multiple messages to the server per request [118, 117, 129]. Fees may be set to 00 in the absence of attack, so that in peaceful times, clients pay nothing [118, 117, 129, 65].
In this paper, we design algorithms for setting the fee (the “price") in an online manner, assisted by an “estimator", which provides statistical information about the distribution of previously occurring good jobs in the stream. Our goal is to set prices so as to minimize the total algorithm’s cost - that is, the cost to the server and honest clients - relative to the attacker’s cost. To achieve this, we define an accuracy measure of the estimator, and determine how the algorithm’s cost increases with this measure.
@@ 1.1 The Dynamic Job Pricing Problem
Server and Clients. The server receives a stream of jobs. Each job is either good: sent by a client - to avoid redundancy, henceforth “clients" rather than “honest clients" - or bad: sent by the adversary (described below).
@@
Estimator. The server has an estimator whose output may be used to set prices. The estimator estimates the number of good jobs generated in any contiguous time interval in the past. For any such interval, $\mathcal{I}$, let $g(\mathcal{I})$ be the number of good jobs in $\mathcal{I}$ and $\hat{g}(\mathcal{I})$ be the value returned by the estimator for interval $\mathcal{I}$. The estimator $\hat{g}$ is additive: for any two non-overlapping intervals, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$:
@@
  • Asynchronous: Communication is asynchronous [87]: there is some unknown maximum latency, $\Delta$ seconds, for any message to be received after it is sent. This latency can also capture the maximum time it took any client to attach a fee. Also, there is some unknown maximum number of good jobs, $M$, that can be generated over $\Delta$ seconds. Both $\Delta$ and $M$ are known to the adversary, but not to the algorithm. Also, the adversary controls the timing of all messages subject to the above constraints.
@@ 1.2 Model Discussion
Rate-limiting is a bandwidth-based defense technique where packets can be dropped with some tunable probability, but this probability is not necessarily shared with the clients [84, 135, 70, 69]; OpenFlow [95, 18, 132, 19, 128] ($>1800$ cites) is a popular such rate-limiting system, even though it has no theoretical guarantees.
@@ 1.3 Our Results
For a given input, let $g$ be the number of good jobs, $B$ be the total spent by the adversary, and $\gamma$ be the estimation gap for the input. Our primary goal is to minimize the algorithmic cost — i.e. fees plus service costs — as a function of $g$, $B$, and $\gamma$
@@
To build intuition for this result, consider the case where $\gamma=O(1)$. Then, the theorem states that the cost to LINEAR is $O(\sqrt{B(g+1)}+(g+1))$. Importantly, the cost to the algorithm grows asymptotically slower than $B$ whenever $g=o(B)$. We note that the additive $O(g+1)$ term results from service costs due to good jobs, which must hold for any algorithm. Our result is actually better than stated in Theorem 1.1, when $\gamma$ grows large; in that case the total cost is always asymptotically bounded by $g^{2}+B$, as will become clear in Section 2.
@@
Moreover, the server sends $O(g\log(B+\gamma))$ messages to clients, and clients sends a total of $O(g\log(B+\gamma))$ messages to the server. Also, each good job is serviced after at most $34\log(\gamma(B+1))+$ 3 messages are sent.
@@
To complement these upper bounds, in Section 0.C, we also prove the following lower bound. This lower bound holds for our zero-latency variant, and so directly also holds for the harder, asynchronous variant. Let $n$ be the total number of jobs.
Theorem 1.3
@@
Finally, we provide simulation results for LINEAR and LINEAR-POWER in our appendix. Our results suggest that the hidden constants in the asymptotic notation may be small, and that the scaling behavior for our algorithms’ costs matches that predicted by theory.
@@ 1.4 Technical Overview
All our proofs are included in the appendix. Here, we give intuition for our technical results. As a warm up, we start with our easiest result: the zero-latency case. Then, we sketch how we build on this result in our asynchronous setting. Finally, we give intuition for our lower bound.
Zero-Latency. Our algorithm for the zero-latency problem is LINEAR. Here, we give a high-level overview of the algorithm and its analysis, see Section 2.1 for detailed pseudocode and Section 0.A for a formal analysis.
@@
Our upper bound consists of four exchange arguments [72], which will be over the ordering of good and bad job requests. Our first two exchange arguments lower bound $B$ and the next two upper bound the algorithm cost, $A$. We begin by lower bounding $B$. To do so, our first exchange argument shows that $B$ is minimized when, in each iteration, the bad jobs occur before the good. Our second exchange argument shows that, subject to this constraint, $B$ is minimized when the bad jobs are distributed as evenly as possible across the iterations. From these two exchange arguments, a lower bound on $B$ follows via an integral lower bound and basic algebra (see Lemma 4).
@@
How do we show that $\delta(b_{i})-\delta(b_{k}-1))\geq 0$? If we relax $b_{i}$ so it can take on values in the real numbers, then $f(b_{i})$ becomes a continuous, differentiable function. Thus, we can use the mean value theorem (MVT). In particular, the MVT shows that $\delta(b_{i})$ equals $f^{\prime}(x)$ at some value of $x\in[b_{i},b_{i+1}]$. Then, upper and lower bounding the $\delta$ values reduces to bounding $f^{\prime}$ in the appropriate range. This is a simpler problem because $f^{\prime}$ is a monotonically non-decreasing (non-increasing) function when $\alpha\geq 1$ ($\alpha<1$, respectively). Once we have this exchange argument in hand, we can calculate the algorithmic costs (see the proofs of Lemma 5 and Theorem 1.1 in our appendix for details).
@@
The first key idea is to partition time into epochs, for the purpose of analysis only. An epoch is a period of time ending with $2\Delta$ seconds during which the current price never increases. Then, several facts follow. First, every good job is serviced within $2$ epochs. Second, we can bound the total fees that all good jobs pay during epoch $i$ to be essentially $O(M(\sqrt{B_{i-1}}+\sqrt{B_{i}}))$, where $B_{j}$ is the amount spent by the adversary in epoch $j$ (see Lemma 11). Third, an epoch has overpaying jobs only if it or its preceding epoch spans the beginning of an iteration (see Lemma 10). This last fact allows us to bound the total number of epochs that add to the overpayment amount using our previous bounds on the number of iterations. Putting all these facts together, we obtain a total cost for LINEAR-POWER that is only $\tilde{O}\big{(}\min\big{(}g+1,M\sqrt{g+1},M\sqrt{B+1}\big{)}\sqrt{B+1}$ $+g\big{)}$ (see Theorem 1.2) versus the cost of LINEAR computed above of $\Theta(\min(g+1,M\sqrt{B+1})\sqrt{B+1}+g)$. Ignoring logarithmic factors, the cost of LINEAR-POWER may be less than LINEAR by a factor of $\sqrt{g+1}/M$. As an example, when $M=O(1)$ and $B=\Theta(g)$, LINEAR-POWER has cost $\tilde{O}(g)$, and LINEAR has cost $\Theta(g^{3/2})$.
Lower Bound. In Section 0.C, we prove a lower bound for the zero-latency case. This bound directly holds for our harder, asynchronous problem variant.
In our lower bound, for given values of $\gamma$, $g_{0}$ and $n$, we design an input distribution that (1) sets $g$ to $g_{0}/\gamma$ with probability $1/2$, and sets $g$ to $\gamma g_{0}$ otherwise; (2) distributes $g$ good jobs uniformly but randomly throughout the input; and (3) creates an estimator that effectively hides both the value of $g$ and also the location of the good jobs, and always has estimation gap at most $\gamma$ on the input (See Section 0.C.1 and Lemma 16 for details).
@@
To address this challenge, we first set the matrix entries to $E(A)-\sqrt{E(B)g\gamma})$, and then prove that the minimax for such a payoff matrix has expected value at least 00. Then, we use Jensen’s inequality and linearity of expectation to show that the minimax is also at least 00 when the matrix entries are $E(A-\sqrt{Bg\gamma})$ (see Lemma 17). Next, using Yao’s principle, we can show the maximin is nonnegative, and thereby obtain a key inequality for $A$ and $B$. Once we have this inequality, we can use algebra to establish our lower bound for any randomized algorithm (see Theorem 1.3).
We believe that this new approach – using Yao’s principle on a matrix whose entries are $X-Y$ in order to bound the ratio $E(X)/E(Y)$, and thereby prove some target lower bound of $E(Y)$ – may have broader applications to proving lower bounds for resource competitive algorithms (see Section 1.6).
1.5 A Roadmap
The remainder of our paper is organized as follows. In Section 1.6, we discuss closely-related work and how it connects with our results. Our algorithms, LINEAR and LINEAR-POWER, are described in Section 2. In Section 3, we give two example estimators, and example runs for both algorithms. We conclude and discuss some future research directions in Section 4. Our appendix includes the upper-bound proofs (Sections 0.A and 0.B), lower-bound argument (Section 0.C), and simulation results (Section 0.D).
@@ 2.1 Algorithm LINEAR
1. If $\hat{g}(\mathcal{I})\geq 1$, where $\mathcal{I}=(t,t^{\prime}]$, $t$ is the start time of the current iteration, and $t^{\prime}$ is the current time, then begin a new iteration. 2. $\textrm{PRICE}\leftarrow s+1$, where $s$ is the number of jobs serviced so far in this iteration. If PRICE has changed, send the new value to all clients. 3. If the next job has a attached fee at least equal to PRICE then service it, otherwise do not service and reply with PRICE. Client: $x\leftarrow 1$.
@@ 2.2 Algorithm LINEAR-POWER
Client. Initially, each client sets a local $x$ variable to $1$. The client sends the server a fee of $x$ with its job. When receiving a message from the server with a $thresh$ value greater than $x$, it sets $x$ to the value received, and re-sends the job with a fee of $x$, or chooses to drop the job.
@@
1. If $\hat{g}(\mathcal{I})\geq 1$, where $\mathcal{I}=(t,t^{\prime}]$, $t$ is the start time of the current iteration, and $t^{\prime}$ is the current time, then begin a new iteration. 2. $\textrm{PRICE}\leftarrow 2^{\left\lfloor\lg(s+1)\right\rfloor}$, where $s$ is the number of jobs serviced in this iteration. 3. If the next job has a fee at least equal to PRICE, then service it. Otherwise, do not service the job; instead send a message with the value PRICE to the client that sent the job. Client: $x\leftarrow 1$.
@@ 3.1 Poisson Estimator
Assume good jobs are generated via a Poisson distribution with parameter $\lambda$, so that the expected interval length between good jobs is $1/\lambda$. Poisson job arrivals are one of the most commonly used statistical models, and are consistent with empirical findings [85, 21, 94, 77].
@@ 3.1.1 Bounding $\gamma$
Hence, the estimation gap on this type of interval is $O(1)$. We conclude that for this example input distribution, w.h.p., $\hat{g}$ is an estimator with $\gamma=O(\log g)$.
@@ 3.2 An Estimator for Weibull, Gamma, and Distributions with Exponentially Bounded Moments
In particular, for any $i\geq 1$, let $X_{i}$ be a random variable giving the time elapsed between the $i-1$ and $i$-th arrival. Assume that the $X_{i}$ values are independent and identically distributed. Then, let $Y_{i}=X_{i}-E(X_{i})$, and assume that for some real positive $L$ and every positive integer $k$,
@@ 3.3 Adversarial Queuing Theory Estimator
To capture a range of network settings, adversarial queuing theory (AQT) [17] parameterizes the behavior of packet arrivals by an injection rate, $\rho$ where $0<\rho\leq 1$, and a non-negative integer, $b$, that denotes burstiness. In AQT, a common arrival model is the leaky-bucket adversary[36] where, for any $t$ consecutive slots, at most $\rho t+b$ packets may arrive, for some number $b$ (for examples, see [50, 49, 6, 31, 5]). The physical metaphor is that there is a bucket of size $b$ gallons. Water enters the bucket according to the (adversarial) arrival process and it leaves at a constant rate of $\rho$. The model requires that $b$ be large enough so that the bucket never overflows.
@@ 3.4 Zero Latency and Asynchronous Example Runs
Zero-Latency. First, suppose the estimator is perfectly accurate. Then, $\hat{g}(\mathcal{I})$ is 00 during an iteration until a good job arrives. Once a good job arrives, $\hat{g}(\mathcal{I})$ increases from 00 to $1$, and a new iteration begins, starting with this new good job. The value PRICE is set back to $1$, and the good job pays a fee of $1$. Then, the cost of each good job is $1$ and the cost of each bad job is the number of bad jobs received since the last good job was received. There are exactly $g$ iterations. If $b$ bad jobs are submitted, the adversary’s costs are minimized when the bad jobs are evenly spread over the iterations, so the cost for each bad job ranges from $1$ to $b/g$ in each of the $g$ iterations, for a total cost of $B=\Theta(b^{2}/g)$.
@@
Figure 3: (Left) LINEAR with iterations of length $1/\lambda$, using the estimator from the zero-latency example in Section 3.4. Black circles are good jobs, red circles are bad. Since the estimation gap is small, each iteration contains about $1$ good job. (Right) Blowup of the last iteration, with $b_{i}$ bad jobs followed by $1$ good job. The adversary’s cost is $B_{i}=\Omega(b^{2}_{i})$ and the algorithm’s cost is $A_{i}=O(b_{i})$, as described in Section 3.4.
@@
Now, assume our example estimator, so that iterations end at the same time as the zero-latency case above. Further, assume a simple distribution for job where exactly $1$ job is generated per iteration at some uniformly distributed time.
Figure 4: The blue circle is a job that is repeatedly bounced, the black circles are other serviced jobs. Arrows are communication from the server to the client of the PRICE value, and from the client to the server of fees. The blue job is bounced $3$ times and pays $1+2+4+8$.
@@
For these jobs $i=2,...,(M-1)$, the adversary can cause the $i$-th job to bounce $\lfloor{\lg(M-i)}\rfloor$ times with payments of $2^{j}$, for $j=0,...,\lfloor{\lg(M-i)}\rfloor$ in the current iteration. Thus, the total cost to the algorithm is $\Omega((g/M)(M^{2}))=\Omega(gM)$.
4 Conclusion and Future Work
Many interesting problems remain. Are there less restrictive properties for the estimator which suffice to achieve a good pricing strategy? Can we model both the clients and the server as selfish but rational agents, as discussed in Section 1.6? In particular, can we design a mechanism that results in a Nash equilibrium, and also provides good cost performance as a function of the adversarial cost?
Appendix
In this appendix, we provide the full proofs for the analysis of LINEAR (Section 0.A) and LINEAR-POWER (Section 0.B), our lower-bound argument (Section 0.C), and our simulation results (Section 0.D).
Appendix 0.A Analysis of LINEAR

### v5

@@ 1 Introduction
It has long been known that fees can be used to curb DoS attacks [45, 63, 114, 96]. This fee may be in the form of computation: a client solves a computational puzzle or Captcha per request [65, 44, 82]; or bandwidth: a client sends multiple messages to the server per request [115, 114, 126]. Fees may be set to 0 in the absence of attack, so that in peaceful times, clients pay nothing.
In this paper, we design algorithms for setting the fee (the “price”) in an online manner, assisted by an “estimator”, which provides statistical information about the distribution of previously occurring good jobs in the stream. Our goal is to set prices so as to minimize the total algorithm’s cost - that is, the cost to the server and honest clients - relative to the attacker’s cost. To achieve this, we define an accuracy measure of the estimator, and determine how the algorithm’s cost increases with this measure.
@@ 1.1 The Dynamic Job Pricing Problem
Server and Clients. The server receives a stream of jobs. Each job is either good: sent by a client - to avoid redundancy, henceforth “clients” rather than “honest clients” - or bad: sent by the adversary (described below).
@@
[…]
@@
[…]
@@ 1.2 Model Discussion
Rate-limiting is a bandwidth-based defense technique where packets can be dropped with some tunable probability, but this probability is not necessarily shared with the clients [83, 132, 70, 69]; OpenFlow [92, 18, 129, 19, 125] (>1800>1800 cites) is a popular such rate-limiting system, even though it has no theoretical guarantees.
@@ 1.3 Our Results
[…]
@@
To build intuition for this result, consider the case where $\gamma=O(1)$. Then, the theorem states that the cost to LINEAR is $O(\sqrt{B(g+1)}+(g+1))$. Importantly, the cost to the algorithm grows asymptotically slower than $B$ whenever $g=o(B)$. We note that the additive $O(g+1)$ term results from service costs due to good jobs, which must hold for any algorithm. Our result is actually better than stated in Theorem 1, when $\gamma$ grows large; in that case the total cost is always bounded by $g^{2}+B$, as will become clear in Section 2.
@@
Moreover, the server sends $O(g\log(B+\gamma))$ messages to clients, and clients sends a total of $O(g\log(B+\gamma))$ messages to the server. Also, each good job is serviced after at most $4\log(\gamma(B+1))+3$ messages are sent.
@@
[…]
Theorem 3.
@@
Finally, in Section 7, we provide preliminary simulation results for LINEAR and LINEAR-POWER. Our results suggest that the hidden constants in the asymptotic notation may be small, and that the scaling behavior for our algorithms’ costs matches that predicted by theory.
@@ 1.4 Technical Overview
Our full proofs are provided in Sections 4, 5, and 6. Here, we give intuition for our technical results. As a warm up, we start with our easiest result: the zero-latency case. Then, we sketch how we build on this result in our asynchronous setting. Finally, we give intuition for our lower bound.
Zero-Latency. Our algorithm for the zero-latency problem is LINEAR. Here, we give a high-level overview of the algorithm and its analysis; our detatiled pseudocode is given in Section 2.1, and our formal analysis is provided in Section 4.
@@
[…]
@@
How do we show that $\delta(b_{i})-\delta(b_{k}-1))\geq 0$? If we relax $b_{i}$ so it can take on values in the real numbers, then $f(b_{i})$ becomes a continuous, differentiable function. Thus, we can use the mean value theorem (MVT). In particular, the MVT shows that $\delta(b_{i})$ equals $f^{\prime}(x)$ at some value of $x\in[b_{i},b_{i+1}]$. Then, upper and lower bounding the $\delta$ values reduces to bounding $f^{\prime}$ in the appropriate range. This is a simpler problem because $f^{\prime}$ is a monotonically non-decreasing (non-increasing) function when $\alpha\geq 1$ ($\alpha<1$, respectively). Once we have this exchange argument in hand, we can calculate the algorithmic costs (see the proofs of Lemma 5 and Theorem 1 in Section 4 for details).
@@
The first key idea is to partition time into epochs, for the purpose of analysis only. An epoch is a period of time ending with $2\Delta$ seconds during which the current price never increases. Then, several facts follow. First, every good job is serviced within $2$ epochs. Second, we can bound the total fees that all good jobs pay during epoch $i$ to be essentially $O(M(\sqrt{B_{i-1}}+\sqrt{B_{i}}))$, where $B_{j}$ is the amount spent by the adversary in epoch $j$ (see Lemma 11). Third, an epoch has overpaying jobs only if it or its preceding epoch spans the beginning of an iteration (see Lemma 10). This last fact allows us to bound the total number of epochs that add to the overpayment amount using our previous bounds on the number of iterations. Putting all these facts together, we obtain a total cost for LINEAR-POWER that is only $\tilde{O}\big(\min\big(g+1,M\sqrt{g+1},M\sqrt{B+1}\big)\sqrt{B+1}$ $g\big)$ (see Section 5) versus the cost of LINEAR computed above of $\Theta(\min(g+1,M\sqrt{B+1})\sqrt{B+1}+g)$. Ignoring logarithmic factors, the cost of LINEAR-POWER may be less than LINEAR by a factor of $\sqrt{g+1}/M$. As an example, when $M=O(1)$ and $B=\Theta(g)$, LINEAR-POWER has cost $\tilde{O}(g)$, and LINEAR has cost $\Theta(g^{3/2})$.
Lower Bound. We prove a lower bound for the zero-latency case (Section 6). This bound directly holds for our harder, asynchronous problem variant.
In our lower bound, for given values of $\gamma$, $g_{0}$ and $n$, we design an input distribution that (1) sets $g$ to $g_{0}/\gamma$ with probability $1/2$, and sets $g$ to $\gamma g_{0}$ otherwise; (2) distributes $g$ good jobs uniformly but randomly throughout the input; and (3) creates an estimator that effectively hides both the value of $g$ and also the location of the good jobs, and always has estimation gap at most $\gamma$ on the input (see Section 6.1 and Lemma 16 for details).
@@
To address this challenge, we first set the matrix entries to $E(A)-\sqrt{E(B)g\gamma})$, and then prove that the minimax for such a payoff matrix has expected value at least 0. Then, we use Jensen’s inequality and linearity of expectation to show that the minimax is also at least 0 when the matrix entries are $E(A-\sqrt{Bg\gamma})$ (see Lemma 17). Next, using Yao’s principle, we can show the maximin is nonnegative, and thereby obtain a key inequality for $A$ and $B$. Once we have this inequality, we can use algebra to establish our lower bound for any randomized algorithm (see Theorem 3).
We believe that this new approach – using Yao’s principle on a matrix whose entries are XYX-Y in order to bound the ratio $E(X)/E(Y)$, and thereby prove some target lower bound of $E(Y)$ – may have broader applications to proving lower bounds for resource competitive algorithms (see our expanded discussion of related work in Section 1.5).
@@ 1.5 Other Related Work
A preliminary version of this work appeared at the 32nd International Colloquium on Structural Information and Communication Complexity (SIROCCO) [24]. Here, we present the full technical analysis, along with an in-depth description and intuition for our results, an expanded discussion of related work, and preliminary findings from our simulations. In addition to the discussion given previously in Section 1.2, we summarize additional related results here.
@@ 2.1 Algorithm LINEAR
1. If $\hat{g}(\mathcal{I})\geq 1$, where $\mathcal{I}=(t,t^{\prime}]$, $t$ is the start time of the current iteration, and $t^{\prime}$ is the current time, then begin a new iteration. 2. $\textrm{PRICE}\leftarrow s+1$, where $s$ is the number of jobs serviced so far in this iteration. If PRICE has changed, send the new value to all clients. 3. If the next job has a attached fee at least equal to PRICE then service it, otherwise do not service and reply with PRICE.
Client: $x\leftarrow 1$.
@@ 2.2 Algorithm LINEAR-POWER
Client. Initially, each client sets a local $x$ variable to $1$. The client sends the server a fee of $x$ with its job. When receiving a message from the server with a threshthresh value greater than $x$, it sets $x$ to the value received, and re-sends the job with a fee of $x$, or chooses to drop the job.
@@
1. If $\hat{g}(\mathcal{I})\geq 1$, where $\mathcal{I}=(t,t^{\prime}]$, $t$ is the start time of the current iteration, and $t^{\prime}$ is the current time, then begin a new iteration. 2. $\textrm{PRICE}\leftarrow 2^{\left\lfloor\lg(s+1)\right\rfloor}$, where $s$ is the number of jobs serviced in this iteration. 3. If the next job has a fee at least equal to PRICE, then service it. Otherwise, do not service the job; instead send a message with the value PRICE to the client that sent the job.
Client: $x\leftarrow 1$.
@@ 3.1 Poisson Estimator
Assume good jobs are generated via a Poisson distribution with parameter $\lambda$, so that the expected interval length between good jobs is $1/\lambda$. Poisson job arrivals are one of the most commonly used statistical models, and are consistent with empirical findings [84, 21].
@@ 3.1.1 Bounding $\gamma$
Hence, the estimation gap on this type of interval is $O(1)$. We conclude that for this example input distribution, w.h.p., $\hat{g}$ is an estimator with $\gamma=O(\log g)$. ∎
@@ 3.2 An Estimator for Weibull, Gamma, and Distributions with Exponentially Bounded Moments
In particular, for any $i\geq 1$, let $X_{i}$ be a random variable giving the time elapsed between the i1i-1 and $i$-th arrival. Assume that the $X_{i}$ values are independent and identically distributed. Then, let $Y_{i}=X_{i}-E(X_{i})$, and assume that for some real positive $L$ and every positive integer $k$,
@@ 3.3 Adversarial Queuing Theory Estimator
[…]
@@ 3.4 Zero Latency and Asynchronous Example Runs
[…]
@@
Figure 3: (Left) LINEAR with iterations of length $1/\lambda$, using the estimator from the zero-latency example in Section 3.4. Black circles are good jobs, red circles are bad. Since the estimation gap is small, each iteration contains about 11 good job. (Right) Blowup of the last iteration, with $b_{i}$ bad jobs followed by $1$ good job. The adversary’s cost is $B_{i}=\Omega(b^{2}_{i})$ and the algorithm’s cost is $A_{i}=O(b_{i})$, as described in Section 3.4.
@@
Now, assume our example estimator, so that iterations end at the same time as the zero-latency case above. Further, assume a simple distribution for jobs where exactly $1$ job is generated per iteration at some uniformly distributed time.
Figure 4: The blue circle is a job that is repeatedly bounced, the black circles are other serviced jobs. Arrows are communication from the server to the client of the PRICE value, and from the client to the server of fees. The blue job is bounced 33 times and pays $1+2+4+8$.
@@
For these jobs i=2,,(M1)i=2,...,(M-1), the adversary can cause the $i$-th job to bounce $\lfloor{\lg(M-i)}\rfloor$ times with payments of $2^{j}$, for $j=0,...,\lfloor{\lg(M-i)}\rfloor$ in the current iteration. Thus, the total cost to the algorithm is $\Omega((g/M)(M^{2}))=\Omega(gM)$.
4 Analysis of LINEAR